OpenStudy (anonymous):

Find t he domain and range: f(x) = x^2 -4x +3 answer D:(-infinity, infinity) R:(-1, infinity)

zepdrix (zepdrix):

\[\large f(x)=x^2-4x+3 \qquad \rightarrow \qquad f(x)=x^2-4x+4-1\] \[\large f(x)=(x-2)^2-1\]-1 is the lower boundary for your range? Hmm yah looks good! :D For some reason your domain bracket looks like an angry smiley to me,\[D:(-\]

OpenStudy (anonymous):

so how do i get domain and range @zepdrix

zepdrix (zepdrix):

Oh i see, you weren't just checking your answer, you need to get it :) my bad. ok one sec.

zepdrix (zepdrix):

@dpaInc u explain whyyyy D:

OpenStudy (anonymous):

lol

zepdrix (zepdrix):

You prolly has been words than my mouf D: with the "rational powers" and uhhh "real coefficients" or blah blah somethign like that D:

zepdrix (zepdrix):

See how I completed the square and put the parabola in standard form fay? :O Does that part make sense?

OpenStudy (anonymous):

i don't see anything wrong... your range is [-1, infinity)

OpenStudy (anonymous):

my graphing calculator just ran out of batteries rite now so -.- yeah..

zepdrix (zepdrix):

She pasted the answer key plank, she wants to know how to get it.

zepdrix (zepdrix):

no batteries!?? so you have to do some ACTUAL work?? :O OH NO!

OpenStudy (anonymous):

yes what u did above i get but how do i find domain and rangestilll

OpenStudy (anonymous):

unless i can find a free online graph. calculator on google?:p

zepdrix (zepdrix):

So remember what a parabola looks like? In standard form, if we were to graph this, we would see that the parabola opens UPWARD because there's no negative in front of the square term. Anddd it's vertex is at the point (2,-1)

OpenStudy (anonymous):

here's the best online GC: https://www.desmos.com/calculator

zepdrix (zepdrix):

Ooooo that looks fun! :O

zepdrix (zepdrix):

can it interpret latex? :X

OpenStudy (anonymous):

wait so just by looking at my graphing calculator i can find domain and range? ^^

OpenStudy (anonymous):

@dpaInc thx man/bro :33